Let X and Y be independent random variables with moment-generating function MX(t) and MY (t) respectively. If a and b are constants and U = aX +bY , show that the moment-generating function of U is MU (t) = MX(at).MY (bt).


1
Expert's answer
2021-11-15T16:03:12-0500

"M_X(t) = E(e^{tX}) \\

M_Y(t) = E(e^{tY})"

X and Y are independent

"U = aX +bY \\

M_U(t) = E(e^{tU}) = E[e^{t(aX+bY)}] \\

 = E[e^{(at)X} \times e^{(bt)Y}] \\

= E[e^{(at)X}] \times E[e^{(bt)Y}] \\

M_U(t) = M_X(at) \times M_Y(bt)"